# Universality in many-body driven systems with an umbilic point

### v1

@@ Acknowledgements
Ž.K. is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). V.P. acknowledges support by ERC Advanced grant No. 101096208 – QUEST, and Deutsche Forschungsgemeinschaft through DFG project KL645/20-2. J.S. and V.P. thank G. Schütz for a discussion.
Appendix A Nonlinear fluctuating hydrodynamics
This Appendix is included for self-containedness. More details can be found in [13]. The dynamics of our many-body stochastic process is a Markov process on an infinite lattice defined via a classical Master equation with rates depicted on Fig.  1. The particle do not change lanes and obey an exclusion rule, i.e. any site can be occupied by at most one particle. The process has two conservation laws, the density of particles $u,v$ on lanes $1,2$, with ranges $10<u,v<$ 1 because of the exclusion rule. Steady states of the process are parametrized by $u,v$, and are factorized over pairs of vertical lattice sites. Namely, denoting by $\eta_{k}^{\alpha}$ a particle number operator on lane $\alpha$ at site $k$, the steady state equal time space correlation satisfy
$\langle\eta_{k}^{\alpha}\eta_{k^{\prime}}^{\beta}\rangle-\langle\eta_{k}^{\alpha}\rangle\langle\eta_{k^{\prime}}^{\beta}\rangle=0,\ \mbox{if }\ k\neq k^{\prime}.$ (10)
@@ Appendix A Nonlinear fluctuating hydrodynamics
$\langle\eta_{k}^{\alpha}\eta_{k}^{\beta}\rangle-\langle\eta_{k}^{\alpha}\rangle\langle\eta_{k}^{\beta}\rangle=C_{\alpha\beta}$ (11)
$C=\left(\begin{array}[]{cc}u(1-u)&-uv+\Omega(u,v)\\ -uv+\Omega(u,v)&v(1-v)\end{array}\right)$ (14)
$\Omega(u,v)=\frac{-1-w_{uv}+\sqrt{(1+w_{uv})^{2}+4(\gamma-1)uv}}{2(\gamma-1)}$ (15)
$w_{uv}=(\gamma-1)(1-u-v)$ (16)
@@
$j_{1}(u,v)=u(1-u)+w_{uv}\Omega+(\gamma-1)\Omega^{2},$ (17)
$j_{2}(u,v)=-j_{1}(v,u).$ (18)
@@
$\partial\vec{\varphi}=-\partial_{x}\left(J\vec{\varphi}+\frac{1}{2}\langle{\varphi}|\vec{H}|{\varphi}\rangle-\partial_{x}\tilde{D}\vec{\varphi}+\tilde{B}\vec{\tilde{\xi}}\right)$ (19)
@@
$J=0$ (20)
$H_{1}=\left(\begin{array}[]{cc}\frac{1-\gamma}{2\sqrt{\gamma}}-2&-\frac{\left(\sqrt{\gamma}-1\right)^{2}}{2\sqrt{\gamma}}\\ -\frac{\left(\sqrt{\gamma}-1\right)^{2}}{2\sqrt{\gamma}}&\frac{1-\gamma}{2\sqrt{\gamma}}\\ \end{array}\right)$ (23)
$H_{2}=-H_{1}-2\sigma^{z}$ (24)
$C=\left(\begin{array}[]{cc}\frac{1}{4}&-\frac{1}{4}+\Omega_{UP}\\ -\frac{1}{4}+\Omega_{UP}&\frac{1}{4}\end{array}\right)$ (27)
$\Omega_{UP}\equiv\Omega\left(\frac{1}{2},\frac{1}{2}\right)=\frac{1}{2+2\sqrt{\gamma}},$ (28)
@@
$R=\left(\begin{array}[]{cc}\frac{1}{\sqrt{2d_{1}}}&0\\ 0&\frac{1}{\sqrt{2d_{2}}}\end{array}\right)\ \left(\begin{array}[]{cc}1&1\\ -1&1\end{array}\right),$ (33)
$d_{1}=\Omega_{UP},\quad d_{2}=\frac{1}{2}-\Omega_{UP}$ (34)
$(RCR^{T})_{\alpha\beta}=\delta_{\alpha\beta}$ (35)
In terms of new variables $\vec{\phi}=R\vec{\varphi}$ at UP Eq. (19) accounting for (20) is rewritten as
$\partial_{t}\vec{\phi}=-\partial_{x}\left(\langle{\phi}|\vec{G}|{\phi}\rangle-\partial_{x}D\vec{\phi}+B\vec{\xi}\right)$ (36)
@@
$G^{\alpha}=\frac{1}{2}\sum_{\beta}R_{\alpha\beta}(R^{-1})^{T}H^{\beta}R^{-1},$ (37)
@@
$G^{1}=b\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$ (40)
$G^{2}=b\left(\begin{array}[]{cc}1&0\\ 0&2-\frac{1}{\sqrt{\gamma}}\end{array}\right)$ (43)
$b=\frac{1}{2}\sqrt{\frac{\sqrt{\gamma}}{\sqrt{\gamma}+1}}$ (44)
Assuming diagonal difusion matrices $D_{\alpha\beta}=D_{\alpha}\delta_{\alpha\beta}$ and writing the system (36) in components, we have
$\partial_{t}\phi_{1}=\partial_{x}(2b\phi_{1}\phi_{2}+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1})$ (45)
$\partial_{t}\phi_{2}=\partial_{x}(b\phi_{1}^{2}+b\left(2-\frac{1}{\sqrt{\gamma}}\right)\phi_{2}^{2}+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2})$ (46)
$\langle\phi_{\alpha}(x),\phi_{\beta}(x^{\prime})\rangle=\delta_{\alpha\beta}\delta(x-x^{\prime})$ (47)
where (47) follows from (35).
The above system of equations, up to phenomenological constants of diffusion and noise terms, coincides with the system of stochastic Burgers equations studied by Spohn and collaborators, see Eq (3.1) in [13].
To access the space-time correlation matrix $S_{\alpha\beta}=\langle\phi(0,0)\phi(x,t)\rangle$ for the system of stochastic Burgers equations (51),(52) we perform a Monte-Carlo study of the two-lane particle system of Fig. 1 for $\gamma=\gamma_{crit}=1/4$. In the Monte-Carlo study we measure the correlation matrix $\tilde{S}_{\alpha\beta}=\langle\varphi(0,0)\varphi(x,t)\rangle$. The change of variables $\phi_{\alpha}=\sum_{\beta=1,2}R_{\alpha\beta}\varphi_{\beta}$ with $R$ from (33) leads to
$S_{11}(x,t)$ $=\frac{1}{2d_{1}}\cdot$ (48)
$\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)+\tilde{S}_{12}(x,t)+\tilde{S}_{21}(x,t)\right)$
$S_{22}(x,t)$ $=\frac{1}{2d_{2}}\cdot$ (49)
$\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)-\tilde{S}_{12}(x,t)-\tilde{S}_{21}(x,t)\right)$
$\frac{1}{2d_{1}}=(1+\gamma^{1/2}),\quad\frac{1}{2d_{2}}=(1+\gamma^{-1/2})$ (50)
reported in the main text, see (5), (6). For $\gamma=\gamma_{crit}=1/4$, the above system acquires an especially simple form
$\partial_{t}\phi_{1}=\partial_{x}(2b\phi_{1}\phi_{2}+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1})$ (51)
$\partial_{t}\phi_{2}=\partial_{x}(b\phi_{1}^{2}+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2})$ (52)
Appendix B Simulation method for two-point functions
@@
$\tilde{\sigma}_{x,t}^{\lambda\mu}(M,\tau,L,R)=$ $\frac{1}{LMR}\sum_{l=1}^{L}\sum_{j=1}^{M}\sum_{r=1}^{R}n_{l+x,j\tau+t}^{\lambda,(r)}n_{l,j\tau}^{\mu,(r)}.$ (53)
@@
$\tilde{S}_{\lambda\mu}(x,t)=\frac{1}{P}\sum_{p=1}^{P}\tilde{\sigma}_{L,x}^{\lambda\mu,(p)}-\rho_{\lambda}\rho_{\mu}+\mathcal{O}(P^{-1/2}).$ (54)
@@
Appendix C Time resolved dynamical exponent
@@
$+(1-p)\int_{0}^{\ln t_{max}}\left|\frac{\mathrm{d}^{2}s_{\alpha}}{\mathrm{d}\tau^{2}}\right|^{2}\mathrm{d}\tau.$ (55)
@@
$S_{\alpha\alpha}(x=0,t)\simeq t^{-2/3}s_{\alpha}(\ln t).$ (56)
@@
$z_{\alpha}(t)=\left(-\frac{\mathrm{d}\ln S_{\alpha\alpha}(0,t)}{\mathrm{d}\ln t}\right)^{-1}$ (57)
@@
$z_{\alpha}(t)\simeq\left(\frac{2}{3}-\frac{1}{s_{\alpha}(\tau)}\left.\frac{\mathrm{d}s_{\alpha}}{\mathrm{d}\tau}\right|_{\ln t}\right)^{-1}.$ (58)

### v2

@@ Acknowledgements
Ž.K. is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). V.P. acknowledges support by ERC Advanced grant No. 101096208 – QUEST, and Deutsche Forschungsgemeinschaft through DFG project KL645/20-2. J.S. and V.P. thank G. Schütz for a discussion. Ž.K. thanks Herbert Spohn for useful discussions. We thank the authors of Ref. [10] for providing their data for the comparison in Figs. D1,D2 of D.
Appendix A
Nonlinear fluctuating hydrodynamics
This Appendix is included for self-containedness. More details can be found in [10]. The dynamics of our many-body stochastic process is a Markov process on an infinite lattice defined via a classical Master equation with rates depicted in Fig. 1. The particle do not change lanes and obey an exclusion rule, i.e. any site can be occupied by at most one particle. The process has two conservation laws, the density of particles $u,v$ on lanes $1,2$, with ranges $10<u,v<$ 1 because of the exclusion rule. Steady states of the process are parametrized by $u,v$, and are factorized over pairs of vertical lattice sites. Namely, denoting by $\eta_{k}^{\alpha}$ a particle number operator on lane $\alpha$ at site $k$, the steady state equal time space correlation satisfy
$\langle\eta_{k}^{\alpha}\eta_{k^{\prime}}^{\beta}\rangle-\langle\eta_{k}^{\alpha}\rangle\langle\eta_{k^{\prime}}^{\beta}\rangle=0,\ \mbox{if }\ k\neq k^{\prime}.$ (A.1)
@@
$\langle\eta_{k}^{\alpha}\eta_{k}^{\beta}\rangle-\langle\eta_{k}^{\alpha}\rangle\langle\eta_{k}^{\beta}\rangle=C_{\alpha\beta}$ (A.2)
$C=\left(\begin{array}[]{cc}u(1-u)&-uv+\Omega(u,v)\\ -uv+\Omega(u,v)&v(1-v)\end{array}\right)$ (A.5)
$\Omega(u,v)=\frac{-1-w_{uv}+\sqrt{(1+w_{uv})^{2}+4(\gamma-1)uv}}{2(\gamma-1)}$ (A.6)
$w_{uv}=(\gamma-1)(1-u-v)$ (A.7)
@@
$j_{1}(u,v)=u(1-u)+w_{uv}\Omega+(\gamma-1)\Omega^{2},$ (A.8)
$j_{2}(u,v)=-j_{1}(v,u).$ (A.9)
@@
$\partial\vec{\varphi}=-\partial_{x}\left(J\vec{\varphi}+\frac{1}{2}\langle{\varphi}|\vec{H}|{\varphi}\rangle-\partial_{x}\tilde{D}\vec{\varphi}+\tilde{B}\vec{\tilde{\xi}}\right)$ (A.10)
@@
$J=0$ (A.11)
$H_{1}=\left(\begin{array}[]{cc}\frac{1-\gamma}{2\gamma^{1/2}}-2&-\frac{\left(\gamma^{1/2}-1\right)^{2}}{2\gamma^{1/2}}\\ -\frac{\left(\gamma^{1/2}-1\right)^{2}}{2\gamma^{1/2}}&\frac{1-\gamma}{2\gamma^{1/2}}\\ \end{array}\right)$ (A.14)
$H_{2}=-H_{1}-2\sigma^{z}$ (A.15)
$C=\left(\begin{array}[]{cc}\frac{1}{4}&-\frac{1}{4}+\Omega_{UP}\\ -\frac{1}{4}+\Omega_{UP}&\frac{1}{4}\end{array}\right)$ (A.18)
$\Omega_{UP}\equiv\Omega\left(\frac{1}{2},\frac{1}{2}\right)=\frac{1}{2+2\gamma^{1/2}},$ (A.19)
@@
$R=\left(\begin{array}[]{cc}\frac{1}{\sqrt{2d_{1}}}&0\\ 0&\frac{1}{\sqrt{2d_{2}}}\end{array}\right)\ \left(\begin{array}[]{cc}1&1\\ -1&1\end{array}\right),$ (A.24)
$d_{1}=\Omega_{UP},\quad d_{2}=\frac{1}{2}-\Omega_{UP}$ (A.25)
$(RCR^{T})_{\alpha\beta}=\delta_{\alpha\beta}$ (A.26)
In terms of new variables $\vec{\phi}=R\vec{\varphi}$ at UP Eq. (A.10) accounting for (A.11) is rewritten as
$\partial_{t}\vec{\phi}=-\partial_{x}\left(\langle{\phi}|\vec{G}|{\phi}\rangle-\partial_{x}D\vec{\phi}+B\vec{\xi}\right)$ (A.27)
@@
$G^{\alpha}=\frac{1}{2}\sum_{\beta}R_{\alpha\beta}(R^{-1})^{T}H^{\beta}R^{-1},$ (A.28)
@@
$G^{1}=b\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)$ (A.31)
$G^{2}=b\left(\begin{array}[]{cc}1&0\\ 0&2-\gamma^{-1/2}\end{array}\right)$ (A.34)
$b=\frac{1}{2}\sqrt{\frac{\gamma^{1/2}}{\gamma^{1/2}+1}}$ (A.35)
Assuming diagonal difusion matrices $D_{\alpha\beta}=D_{\alpha}\delta_{\alpha\beta}$ and writing the system (A.27) in components, we have
$\partial_{t}\phi_{1}=\partial_{x}(2b\phi_{1}\phi_{2}+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1})$ (A.36)
$\partial_{t}\phi_{2}=\partial_{x}(b\phi_{1}^{2}+b\left(2-\gamma^{-1/2}\right)\phi_{2}^{2}+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2})$ (A.37)
$\langle\phi_{\alpha}(x),\phi_{\beta}(x^{\prime})\rangle=\delta_{\alpha\beta}\delta(x-x^{\prime})$ (A.38)
where (A.38) follows from (A.26). The above system of equations, up to phenomenological constants of diffusion and noise terms, coincides with the system of stochastic Burgers equations studied by Spohn and collaborators, see Eq (3.1) in [10].
Note that the system of Eqs. (A.36,A.37) is not invariant under the Galilean transformation
$\phi_{2}(x,t)=B+\phi_{2}(x+2bBt,t)$ (A.39)
$\phi_{1}(x,t)=\phi_{1}(x+2bBt,t),$ (A.40)
except at the point $12-\gamma^{-1/2}=$ 1, or $\gamma=1$ i.e. at the point where the system splits into the two independent one-lane exclusion processes, governed by the KPZ universality, see Fig. 1 and Fig. 6. Away from the point $\gamma=1$ the absence of Galilean invariance, as argued in [18, 19, 20], breaks the validity of the scaling identity $\xi_{2}+z_{2}=2$ between the roughening exponent $\xi_{2}$ (equal to $\frac{1}{2}$ ) and the dynamical exponent $z_{2}$, yielding the dynamical exponents $z_{2}$ (and similarly, $z_{1}$) free of constraints.
To access the space-time correlation matrix $S_{\alpha\beta}=\langle\phi(0,0)\phi(x,t)\rangle$ for the system of stochastic Burgers equations (A.44),(A.45) we perform a Monte-Carlo study of the two-lane particle system of Fig. 1 for $\gamma=\gamma_{crit}=1/4$. In the Monte-Carlo study we measure the correlation matrix $\tilde{S}_{\alpha\beta}=\langle\varphi(0,0)\varphi(x,t)\rangle$. The change of variables $\phi_{\alpha}=\sum_{\beta=1,2}R_{\alpha\beta}\varphi_{\beta}$ with $R$ from (A.24) leads to
$S_{11}(x,t)$ $=\frac{1}{2d_{1}}\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)+\tilde{S}_{12}(x,t)+\tilde{S}_{21}(x,t)\right)$ (A.41)
$S_{22}(x,t)$ $=\frac{1}{2d_{2}}\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)-\tilde{S}_{12}(x,t)-\tilde{S}_{21}(x,t)\right)$ (A.42)
$\frac{1}{2d_{1}}=(1+\gamma^{1/2}),\quad\frac{1}{2d_{2}}=(1+\gamma^{-1/2})$ (A.43)
reported in the main text, see (3.2), (3.3). For $\gamma=\gamma_{crit}=1/4$, the above system acquires an especially simple form
$\partial_{t}\phi_{1}=\partial_{x}(2b\phi_{1}\phi_{2}+D_{1}\partial_{x}\phi_{1}+B_{1}\xi_{1})$ (A.44)
$\partial_{t}\phi_{2}=\partial_{x}(b\phi_{1}^{2}+D_{2}\partial_{x}\phi_{2}+B_{2}\xi_{2})$ (A.45)
Appendix B
Simulation method for two-point functions
@@
$\tilde{\sigma}_{x,t}^{\lambda\mu}(M,\tau,L,R)=$ $\frac{1}{LMR}\sum_{l=1}^{L}\sum_{j=1}^{M}\sum_{r=1}^{R}n_{l+x,j\tau+t}^{\lambda,(r)}n_{l,j\tau}^{\mu,(r)}.$ (B.1)
@@
$\tilde{S}_{\lambda\mu}(x,t)=\frac{1}{P}\sum_{p=1}^{P}\tilde{\sigma}_{L,x}^{\lambda\mu,(p)}-\rho_{\lambda}\rho_{\mu}+\mathcal{O}(P^{-1/2}).$ (B.2)
@@
Appendix C
Time resolved dynamical exponent
@@
$+(1-p)\int_{0}^{\ln t_{max}}\left|\frac{\mathrm{d}^{2}s_{\alpha}}{\mathrm{d}\tau^{2}}\right|^{2}\mathrm{d}\tau.$ (C.1)
@@
$S_{\alpha\alpha}(x=0,t)\simeq t^{-2/3}s_{\alpha}(\ln t).$ (C.2)
@@
$z_{\alpha}(t)=\left(-\frac{\mathrm{d}\ln S_{\alpha\alpha}(0,t)}{\mathrm{d}\ln t}\right)^{-1}$ (C.3)
@@
$z_{\alpha}(t)\simeq\left(\frac{2}{3}-\frac{1}{s_{\alpha}(\tau)}\left.\frac{\mathrm{d}s_{\alpha}}{\mathrm{d}\tau}\right|_{\ln t}\right)^{-1}.$ (C.4)
Appendix D
Comparison of intermediate-time data with the data of Ref. [10]
In this Appendix we compare our data against the data of Figs.1,4 of Ref. [10]. In Figs. 8 and 9 we show the comparisons of space-time correlators of the two-lane lattice model and of the system of coupled Burgers equations respectively. In both cases our data (shown by colored lines) agree with the data of Ref. [10] at comparable times. We note that while the dynamical exponent $z_{\rm KPZ}=3/2$ appears to fit the data rather well (only the off-diagonal correlator in the lattice models fails to collapse fully), this is only due to the insufficiently short time window (the maximal and the minimal times in Fig.1 and in Fig.4 of Ref. [10] differ just by a factor of 2 ). To observe the deviations from $z_{\rm KPZ}=3/2$ a sufficiently long time window is needed, e.g. in our Fig. 5 we have $t_{\max}/t_{\min}>300)$.
Figure 8: Comparison of structure factors of the two-lane Markov process (colored lines) scaled with dynamical exponent $z=3/2$ against data presented in Figure 4 of Ref. [10] at comparable times (colored crosses) . Simulation parameters as in Fig. 3.
Figure 9: Comparison of structure factors of the system of coupled Burgers equations (3.6),(3.7) (colored lines) scaled with dynamical exponent $z_{\rm KPZ}=3/2$ against data in Figure 1 of Ref. [10] at comparable times (black squares, averaged over all three times $t=200,300,400$) . Simulation parameters as in Fig. 7.